A metal sphere with a negative charge of 3.00 µC is placed 12.0 cm from another similar metal sphere with a positive charge of 2.00 µC. The two spheres momentarily touch, and then return to their original positions. Calculate the electrostatic force acting on the two metal spheres. 


1
Expert's answer
2021-10-05T10:05:45-0400

As the two spheres are identical then after contact charge in them will be equal and will be half of the total charge.

Charge of the first sphere is

"Q_1=-3\u00b5C"

charge of the second sphere is

"Q_2=2\u00b5C"

After contact charge in each sphere is

"Q=\frac{Q_1+Q_2}{2} \\

Q=\frac{-3µ C+2µC}{2}=-0.5µ C=-0.5\times 10^{-6}C"

Distance between them is

d=12cm =0.12 m

So now electrostatic force between them is

"F=k\frac{Q^2}{d^2} \\

Wherek=9\times 10^9Nm^2/C^2 \\

\therefore F=9\times 10^9\times \frac{(0.5\times10^{-6})^2}{(0.12)^2} \\

F=1.56\times 10^{-1}N"

As nature of the two charged sphere are same.


Therefore the force between them is repulsive